The overall reaction in a commercial heat pack can
Answer #1

a)

n = 4.2 mol of Fe react

-1652 kJ accoridng to the Hrxn based on 4 mol

b)

when 1 mol is produced,

since the Hrxn is based in 2 moles:

1/2*Hrxn = 1/2*-1652 = -826 kJ

c)

1 g of Fe

mol = masS/MW = 1/55.5 = 0.0180

Q = n*Hrxn = 0.0180*(-1652/4) = -7.434 kJ

d)

12.5 g of Fe burn

mol = 12.5/55.5 = 0.22727

Q = n*Hrxn = 0.22727*(-1652/4) = -93.86251 kJ